4 out of 12 television set chosen at of random are defentive, if three of the tv set are choosen at of random for shipment to a hotel, how many defective set can they expect


1
Expert's answer
2021-02-26T00:53:15-0500

Probability of choosing defective televisions is (p)"= \dfrac{4}{12}= \dfrac{1}{3}"


Number of televisions that are choose for shipment to a hotel(n) = 3


So, the number of defective set they can expect= "n\times p=3\times \dfrac{1}{3}=1"


Hence they should expect 1 defective television.